OpenStudy (anonymous):

A rubber ball dropped from a height of 1250 inches. After each bounce it rebounds 2 fifths of the distance it fell. How far does the ball fall on it's fourth fall?

OpenStudy (jhannybean):

@dan815

OpenStudy (anonymous):

I believe it is 500 inches.

OpenStudy (anonymous):

This was my equation on solving it: x = 1250 inches x (2/5 x 4) x = 1250 x (8/20) x = 10000/20 x = 500

OpenStudy (anonymous):

Am I right? @Jhannybean